brando17:

If AF = 23 cm, Find the length of arc EC

mdobbs6856:

Do you have a picture?

brando17:

yes

brando17:

mdobbs6856:

@dude

surjithayer:

arc EC=180-87=93 AF=23 cm radius r=23 cm \[\theta=93^\circ =\frac{ \pi }{ 180 } \times 93=\frac{ 93 \pi }{ 180}~radians\] \[l=r \theta=23 \times \frac{ 93 \pi }{ 180} \approx 37.33 cm\]

Extrinix:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @surjithayer arc EC=180-87=93 AF=23 cm radius r=23 cm \[\theta=93^\circ =\frac{ \pi }{ 180 } \times 93=\frac{ 93 \pi }{ 180}~radians\] \[l=r \theta=23 \times \frac{ 93 \pi }{ 180} \approx 37.33 cm\] \(\color{#0cbb34}{\text{End of Quote}}\) (next time fyi \(explain\) how you got to the answer and \(do not\) give a direct answer, its against te rules for both)